# Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning

### v2

@@ 1 Introduction
In recent years, with the rapid development of large language models (LLMs), Reinforcement learning (RL) has once again become a key technique for post-training and human alignment. In particular, DeepSeek-R1 [guo2025deepseek] introduced the Group Relative Policy Optimization (GRPO) algorithm, which successfully incentivized LLMs to exhibit stronger reasoning abilities in complex tasks. This work highlighted the potential of RL in improving reasoning and alignment in large models, and it has further stimulated widespread interest in efficient RL training methods across both academia and industry.
Nevertheless, despite the remarkable potential of RL in LLM alignment and reasoning tasks, training efficiency remains a critical challenge. This stems from the inherent complexity of the RL training pipeline. First, even the forward computation alone involves the primary training model (policy), an old policy model, and a reference model, while certain samples may additionally require a value model and a reward model [schulman2017proximal], all of which substantially increase computational overhead. Second, before each training step, the latest weights of the policy model are often used to generate a large number of chain-of-thought (CoT) [wei2022chain] trajectories to support optimization, further amplifying the inference burden. These factors not only constrain training throughput but also result in significant memory consumption, making efficiency improvements particularly challenging in large-scale model settings.
The optimization of reinforcement learning training has evolved continuously over time. Early DeepSpeed-Chat [yao2023dschat] primarily relied on ZeRO for distributed training, where the policy model itself was directly responsible for inference. This approach suffered from relatively low inference efficiency, which limited overall training throughput. Subsequently, an architecture that decouples inference from training is introduced, utilizing specialized inference engines such as vLLM [kwon2023efficient] to generate responses. Although these approaches increased resource usage slightly, they significantly improved inference efficiency, thereby substantially enhancing overall training performance, as exemplified by Open-R1 [openr1]. To further leverage computing resources and reduce idle time between training and inference, the industry introduced training–inference co-location schemes. These approaches are conceptually still based on training–inference separation but deploy both tasks within the same resource pool, sharing resources through repeated task switching. Typical instances include OpenRLHF [hu2024openrlhf]. However, such architecture remain tightly coupled, limiting independent scalability and preventing training and inference from executing simultaneously. Beyond optimizing inference, additional exploration has focused on the training frameworks themselves. For example, MindSpeed-RL [feng2025mindspeedrldistributeddataflow] and VERL [sheng2024hybridflow] have adopted Megatron-style 3D parallelism [megatron-lm] to train the policy model, improving throughput and enabling the training of larger models.
Against this backdrop, we revisit the inference–training separation strategy and enhance a conventional synchronous framework with new optimizations. Our approach builds on a Megatron-style 3D parallel architecture, further refined for improved training efficiency. Specifically, we add a background process in the data-loading module that acts as a producer, responsible for sending a full batch of prompts to the inferrence service. The main process acts as a consumer, immediately retrieving the outputs for training, where the received samples belonging to the same group are further divided into multiple micro-batches and trained sequentially. Once the entire batch has been consumed, the policy model’s weights are synchronized to the rollout workers, after which the next batch of asynchronous training begins. We refer to this mechanism as periodic asynchrony, which remains a purely on-policy asynchronous strategy and can thus be applied to any on-policy RL algorithm without modification. Benefiting from the decoupling of inference and training, the number of inference instances can be flexibly scaled according to inference efficiency, preventing inference from becoming the bottleneck. Furthermore, in GRPO, where each group of samples shares the same prompt, we propose a shared-prompt attention mask to reduce redundant computation and further improve efficiency. Experimental results demonstrate that, using 16 NPUs to train an 8B dense model with a 16k context length, our method achieves more than a threefold improvement in end-to-end RL training throughput.
@@ 2 Related Work
Improving the efficiency of reinforcement learning (RL) for large language models (LLMs) has been the subject of growing attention in both academia and industry. While synchronous frameworks dominate existing large-scale RL systems due to their algorithmic stability, they often suffer from underutilization of computational resources. To overcome these limitations, recent studies have explored asynchronous training paradigms, which aim to decouple inference from training and thereby enhance throughput and scalability. In this section, we review representative efforts in this direction, with a particular emphasis on asynchronous reinforcement learning approaches.
@@
One of the earliest attempts in this direction is Asynchronous RLHF [noukhovitch2024asynchronous], which decouples the generation and learning processes, allowing new sample generation and old sample training to proceed in parallel. Its core idea is online but off-policy RLHF, where the model is updated using samples from previous iterations. While effective, this represents a typical off-policy strategy that may not directly apply to on-policy algorithms such as GRPO, and it remains constrained to relatively short-context settings.
@@
Most recently, AReaL [fu2025areallargescaleasynchronousreinforcement] advanced the application of asynchronous RL to LLM reasoning tasks. After we had already developed the methods presented in this paper, we became aware of AReaL, which fully decouples generation from training while introducing a parameter $\eta$ to control data staleness. To better handle delayed samples, it incorporates a staleness-enhanced PPO variant, effectively modifying the on-policy paradigm to tolerate partially off-policy data. Additionally, its training start time depends on the completion of the first inference batch. Similarly, during the preparation of this manuscript, ROLL Flash [lu2025iirollflash] introduced a study sharing conceptual similarities with AReaL’s work. It also decouples training from inference and controls the proportion of off-policy data through an async ratio. Both AReaL and ROLL Flash are off-policy asynchronous frameworks that improve training efficiency but require modifications to standard on-policy RL algorithms to handle data bias. This design trades a small portion of data originally intended for strictly on-policy training to avoid inference-service blocking, effectively shifting it toward off-policy learning. Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
However, they all fall under off-policy asynchronous frameworks, requiring control of data bias and modifications to algorithms to maintain stability for on-policy model training. Distinct from these approaches, our method introduces a new perspective with several key innovations:
  1. 1.
@@
    Periodic asynchrony, which guarantees purely on-policy asynchronous acceleration without compromising algorithmic stability.
  2. 2.
@@
    Decoupling of prompt-level inference production and sample-level training consumption, which maximally overlaps inference and training.
  3. 3.
    A group-mask design that enables effective reuse of forward and backward computations for prompt tokens, thereby reducing redundant computation.
  4. 4.
@@
    A distribution-consistent design across the policy model, old-policy model, and reference model, allowing more flexible scaling and simplifying the overall system.
  5. 5.
@@
    Independent scalability of inference and training instances, ensuring that neither component becomes a performance bottleneck.
@@
Together, these designs enable efficient long-context training with substantially fewer computational resources. All the components have been implemented in the project release v5.0.5 on September 5, 2025 (https://github.com/janelu9/EasyLLM).
@@ 3 Background
We can formally demonstrate the equivalence of full-batch and micro-batch training by reviewing the gradient update rules. When training on an entire batch directly, the model parameters are updated as follows:
$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n}$ (1)
where $N$ is the total number of samples in the batch, and $g_{n}$ denotes the gradient computed from the $n$-th sample.
For micro-batch training, the update proceeds differently. First, the accumulated gradient is initialized. Then, for each micro-batch $j=1$ to $M$, the gradient is updated as
$g_{\text{acc}}:=g_{\text{acc}}+\frac{1}{m}\sum_{i=1}^{m}g_{i,j},$ (2)
where $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch. Finally, after processing all micro-batches, the model parameters are updated using the accumulated gradient:
$W_{t+1}=W_{t}-\alpha\frac{1}{M}g_{\text{acc}}.$ (3)
This formulation demonstrates that, mathematically, full-batch training and micro-batch training are equivalent. However, the micro-batch approach requires significantly less GPU memory. In particular, this delayed parameter update scheme is more suitable for reinforcement learning.
@@
$J_{GRPO}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{old},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{ref})\Big),$ (4)
where
$L_{i}(\pi_{\theta}\|\pi_{old},A_{i})=\min\Bigg(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},\ \text{clip}\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\Big)A_{i}\Bigg),$ (5)
$D_{KL}^{i}(\pi_{\theta}\|\pi_{ref})=\frac{\pi_{ref}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\frac{\pi_{ref}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (6)
For a batch of $N$ prompts, the objective function becomes:
$J_{batch}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{GRPO}^{i}$ (7)
@@
Then, the batch can be trained using micro-batches as:
$J_{batch}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (9)
@@ 4.1 Unified Tri-Model Architecture
    Synchronize the weights of the policy model to the inference engine.
@@
    The inference engine retrieves prompts from the dataloader and generates responses. A reward module evaluates the generated responses based on the prompts, distinguishing high-quality responses from lower-quality ones.
@@
    All generated samples and their associated scores are sent to the training engine for loss computation and backpropagation, updating the weights of the policy model.
@@
Figure 1: Architecture of unified triple models with shared distribution.
@@ 4.2 Shared-Prompt Attention
Compared with standard training, when adopting the shared-prompt approach, the following adjustments are required:
@@
    Concatenate the shared prompt with the token IDs of multiple responses. Correspondingly, the labels used for computing probabilities should also be concatenated, except that the prompt portion does not need to be included in the labels.
@@
    The position indices need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
    $p=\left({0,1,...,\left|{{x_{p}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{1}}}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{2}}}}}\right|-1}\right)$ (11)
@@
    In the attention module, use a group mask as illustrated in the Figure 2, instead of the standard causal mask. This ensures that the value vectors of the second and subsequent responses are only aggregated with the prompt’s value vectors, and not with the value vectors of the preceding responses.
@@
    $\left\{\begin{array}[]{l}\hat{y}={\mathop{\rm model}\nolimits}\left({x,mas{k_{group}}}\right)\\ {{\hat{y}}^{old}}={{\mathop{\rm model}\nolimits}^{old}}\left({x,mas{k_{group}}}\right)\\ {{\hat{y}}^{ref}}={{\mathop{\rm model}\nolimits}^{ref}}\left({x,mas{k_{group}}}\right)\end{array}\right$. (12)
@@
    When computing probabilities, truncate the prompt tokens and retain only the response portion. Finally, the probabilities output by the three types of models, together with the reward scores, are fed into the GRPO objective function to compute the loss and perform iterative training.
    $\left\{\begin{array}[]{l}\pi=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}_{\left|{{x_{p}}}\right|:\left|x\right|}},y}\right)\\ {\pi_{old}}=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}^{old}}_{\left|{{x_{p}}}\right|:\left|x\right|},y}\right)\\ {\pi_{ref}}=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}^{ref}}_{\left|{{x_{p}}}\right|:\left|x\right|},y}\right)\\ Loss={J_{GRPO}}\left({\pi,{\pi_{old}},{\pi_{ref}},A}\right)\end{array}\right$. (13)
@@ 4.3 Periodic Asynchronous Reinforcement Learning
Asynchronous execution of inference and training is a key factor in accelerating reinforcement learning systems with a separated training-inference architecture. In such asynchronous systems, the critical aspect is minimizing the time the training engine waits for the rollout workers to return generated samples, i.e., the startup latency. In this work, we propose a simple approach that introduces a temporary data generator between the data loader and the trainer. This design transforms on-policy reinforcement learning from synchronous to asynchronous execution, significantly improving training efficiency without requiring any modifications to the underlying reinforcement learning algorithm. The reinforcement learning workflow incorporating the temporary sample generator is illustrated in Figure 3. It follows a typical producer–consumer pattern and consists of four main components:
@@
Figure 3: How our asynchronous system operates.
@@
1:Dataloader $D$, training iterations $T$, batch size $B$
2:Trained policy model
3:Initialize shared queue $Q$ and gradient accumulator $O$
4:for $t=1$ to $T$ do
5:  Synchronize policy model weights to the rollout workers
6:  Retrieve prompts $P=\{p_{1},\dots,p_{B}\}$ from $D$
7:  Launch thread Producer:
8:   Create $B$ concurrent coroutine tasks
9:   Run in parallel for each $i=1,\dots,B$:
10:    $r_{i}\leftarrow$ InferenceService($p_{i}$)
11:    $a_{i}\leftarrow$ RewardFunction($r_{i}$)
12:    Enqueue $(a_{i},r_{i})$ into $Q$
13:  $O\leftarrow 0$
14:  $consumed\_count\leftarrow 0$
15:  while consumed_count<Bconsumed\_count<B do
16:     $(a_{i},r_{i})\leftarrow$ Dequeue($Q$)
17:     $s_{i}\leftarrow$ ProcessSample($p_{i},r_{i}$)
18:     Compute loss $L(s_{i})$ then backward
19:     $O\leftarrow O+\nabla L(s_{i})$
20:     $consumed\_count\leftarrow consumed\_count+1$
21:  end while
22:  Move policy’s weights into old-policy’s state_dict
23:  Update policy’s weights using gradient $O$
24:end for
(a) Training one step in a synchronous system.
(b) Training two steps in our asynchronous system.
Figure 4: Comparison of training steps in synchronous and our asynchronous systems.

### v3

@@ 1 Introduction
Reinforcement learning (RL) has re-emerged as a key technique for post-training and aligning large language models (LLMs). Following the introduction of GRPO by DeepSeek-R1 Guo et al. (2025), RL has demonstrated strong potential in improving reasoning capabilities, sparking growing interest in efficient RL pipelines across academia and industry.
Despite these advances, RL for LLMs still faces severe efficiency challenges. The training pipeline requires multiple models in the forward pass—the policy model, old-policy model, and reference model, and in some cases a value and reward model Schulman et al. (2017)—leading to substantial computational overhead. Moreover, each training step typically depends on generating large numbers of chain-of-thought (CoT) Wei et al. (2022) trajectories using the latest policy weights, further increasing inference cost and memory usage.
To improve efficiency, early systems such as DeepSpeed-Chat Yao et al. (2023) relied on ZeRO-based distributed training, where the policy model handled inference directly, resulting in limited throughput. Later approaches decoupled training and inference, leveraging inference engines like vLLM Kwon et al. (2023) to accelerate rollout generation, as demonstrated by Open-R1 Hugging Face (2025). Industry systems such as OpenRLHF Hu et al. (2024) push this further through training–inference co-location, although task switching still prevents full concurrency and independent scalability. Meanwhile, frameworks such as MindSpeed-RL Feng et al. (2025) and VERL Sheng et al. (2024) improve the training side through Megatron-style 3D parallelism Shoeybi et al. (2019).
Motivated by these limitations, we revisit the training–inference separation strategy and introduce a synchronous framework enhanced with new system-level optimizations. Our design uses a Megatron-style 3D parallel backbone and adds a background producer process that continuously dispatches prompts to inference workers, while the main consumer process retrieves outputs and performs training. Samples from each group are split into micro-batches and trained sequentially, and model weights are synchronized only after the full batch is consumed. This periodic asynchrony achieves purely on-policy asynchronous acceleration and is compatible with any on-policy RL algorithm. The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation. Experiments show that on a 16-NPU setup training an 8B model with 16k context, our approach delivers over a threefold improvement in end-to-end RL throughput.
@@ 2 Related Work
Improving the efficiency of reinforcement learning (RL) for large language models (LLMs) has attracted increasing attention across academia and industry. Although synchronous frameworks dominate large-scale RL systems due to their stability, they often underutilize computational resources. To address this, recent studies explore asynchronous training paradigms that decouple inference from training to enhance throughput and scalability. We review representative efforts, with emphasis on asynchronous RL methods.
@@
An early attempt is Asynchronous RLHF Noukhovitch et al. (2024), which decouples generation and learning so that new sample generation and old-sample training proceed in parallel. Its key idea is online but off-policy RLHF, updating models with samples from previous iterations. While effective, it remains a typical off-policy strategy and is less suitable for strictly on-policy algorithms such as GRPO, especially in long-context settings.
@@
Most recently, AReaL Fu et al. (2025) advanced the application of asynchronous RL to LLM reasoning tasks by fully decoupling generation from training. It introduces a parameter $\eta$ to control data staleness, enabling better handling of delayed samples. To better handle delayed samples, it incorporates a staleness-enhanced PPO variant, effectively modifying the on-policy paradigm to tolerate partially off-policy data. Additionally, its training start time depends on the completion of the first inference batch. In a similar vein, ROLL Flash Lu et al. (2025), also recently introduced, shares conceptual similarities with AReaL. Like AReaL, it decouples training from inference and controls the proportion of off-policy data through an async ratio. Both AReaL and ROLL Flash are off-policy asynchronous frameworks that improve training efficiency but require modifications to standard on-policy RL algorithms to handle data bias. This design trades a small portion of data originally intended for strictly on-policy training to avoid inference-service blocking, effectively shifting it toward off-policy learning. Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
In contrast, our method takes a different perspective with several innovations:
@@
  • Algorithm-decoupled periodic asynchrony enabling purely on-policy asynchronous acceleration without sacrificing stability.
@@
  • Decoupling prompt-level inference production from sample-level training consumption, maximizing overlap between the two.
  • A shared-prompt mask for efficient reuse of forward/backward computation on prompt tokens.
@@
  • A distribution-consistent design across policy, old-policy, and reference models, simplifying scaling and system structure.
@@
  • Independent scalability of inference and training instances, preventing bottlenecks in either component.
@@
Together, these designs enable efficient long-context training with substantially fewer computational resources. All components are implemented in project release v5.0.5 (September 5, 2025): https://github.com/janelu9/EasyLLM.
@@ 3 Background
The gradient update rule for full-batch training is:
$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n},$ (1)
and for micro-batch training, after accumulating gradients over all micro-batches, the update rule is:
$W_{t+1}=W_{t}-\alpha\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}g_{i,j},$ (2)
where $M$ is the total number of micro-batches, $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch. This demonstrates that both training strategies are mathematically equivalent, but micro-batching offers significant memory savings.
@@
$J_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big).$ (3)
The PPO-style clipped advantage term can be split into two lines as follows:
$L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})=\min\Bigg($ $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)}A_{i},$
$\text{clip}\!\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)},1-\varepsilon,\,1+\varepsilon\Big)A_{i}\Bigg).$ (4)
The KL divergence term is given by:
$D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})=\frac{\pi_{\mathrm{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\frac{\pi_{\mathrm{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (5)
For a batch of $N$ prompts, the objective becomes:
$J_{\mathrm{batch}}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{\mathrm{GRPO}}^{i}$
@@
Then the batch objective can be rewritten using micro-batches:
$J_{\mathrm{batch}}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (8)
@@ 4.1 Unified Tri-Model Architecture
    Synchronize the weights of the policy model with the inference engine.
@@
    The inference engine retrieves prompts from the dataloader and generates responses. A reward module evaluates these responses based on the corresponding prompts, distinguishing high-quality outputs from lower-quality ones.
@@
    The generated samples and their associated scores are then sent to the training engine, which computes the loss and performs backpropagation to update the policy model parameters.
@@
Figure 1: Triple Models with Shared Distribution.
@@ 4.2 Shared-Prompt Attention
Compared with standard training, when adopting the shared-prompt approach, the following modifications are made:
@@
    Concatenate the shared prompt with the token IDs of multiple responses. The corresponding labels are concatenated accordingly, while the prompt portion is excluded from the labels.
@@
    In this formulation, $x_{p}$ denotes the token IDs of the shared prompt, and $x_{r_{1}},x_{r_{2}}$ correspond to the tokens of the individual responses.
@@
    Rearrange the position indices such that each response starts immediately after the prompt, ensuring positional continuity with the shared prompt.
    $p=($ $0,1,\dots,|x_{p}|-1,\ |x_{p}|,\dots,|x_{p}|+|x_{r_{1}}|-1,$
    $|x_{p}|,\dots,|x_{p}|+|x_{r_{2}}|-1)$ (10)
    Here, the notation $|\cdot|$ indicates sequence length.
@@
    In the attention module, a shared-prompt mask (Figure 2) is used instead of the standard causal mask. This design restricts each response token to attend only to the shared prompt, preventing information leakage across responses.
@@
    For probability computation, the prompt tokens are discarded and the loss is computed only over the response tokens:
    $\pi=-\mathop{\rm CrossEntropy}\nolimits\left(\hat{y}_{|x_{p}|:|x|},y\right),$ (11)
    where $\hat{y}_{|x_{p}|:|x|}$ represents the predicted logits of the response tokens.
@@ 4.3 Periodic Asynchronous Reinforcement Learning
Asynchronous execution of inference and training is a key factor in accelerating reinforcement learning systems with a separated training-inference architecture. In such asynchronous systems, the critical aspect is minimizing the time the training engine waits for the rollout workers to return generated samples, i.e., the startup latency. In this work, we propose a simple approach that introduces a temporary data generator between the data loader and the trainer. This design transforms on-policy reinforcement learning from synchronous to asynchronous execution, significantly improving training efficiency without requiring any modifications to the underlying reinforcement learning algorithm.
The reinforcement learning workflow incorporating the temporary sample generator is illustrated in Figure 3. It follows a typical producer–consumer pattern and consists of four main components:
@@
Figure 3: Operational Mechanism of the Asynchronous System.
@@
Input: Dataloader $D$, training iterations $T$, batch size $B$
Output: Trained policy model
1: Initialize shared queue $Q$ and gradient accumulator $O$
2:for $t=1$ to $T$ do
3:  Synchronize policy model weights to the rollout workers
4:  Retrieve prompts $P=\{p_{1},\dots,p_{B}\}$ from $D$
5:  Producer thread:
6:     Create $B$ concurrent coroutine tasks
7:     Run in parallel for each $i=1,\dots,B$:
8:      $r_{i}\leftarrow$ InferenceService($p_{i}$)
9:      $a_{i}\leftarrow$ RewardFunction($r_{i}$)
10:      Enqueue $(a_{i},r_{i})$ into $Q$
11:  $O\leftarrow 0$
12:  $consumed\_count\leftarrow 0$
13:  while consumed_count<Bconsumed\_count<B do
14:     $(a_{i},r_{i})\leftarrow$ Dequeue($Q$)
15:     $s_{i}\leftarrow$ ProcessSample($p_{i},r_{i}$)
16:     Compute loss $L(s_{i})$ then backward
17:     $O\leftarrow O+\nabla L(s_{i})$
18:     $consumed\_count\leftarrow consumed\_count+1$
19:  end while
20:  Move policy’s weights into old-policy’s modules
21:  Update policy’s weights using gradient $O$
22:end for
@@
(a) Synchronous training step.
(b) Asynchronous two-step training.
Figure 4: Comparison of synchronous vs. asynchronous training.